12) Suppose that a room contains 10 cats and 15 dogs. How many ways are there to form a committee consisting of 6 of these animals if it must have more dogs than cats?
1
Expert's answer
2020-11-17T06:06:48-0500

This committe must consist of 6 dogs or 5 dogs and 1 cat or 4 dogs and 2 cats. The number of ways of choosing "k" elements from "n" elements is equal to "\left(\begin{array}{c}n\\k\end{array}\right)=\frac{n!}{k!(n-k)!}"

Therefore, using combinatorial Sum and Product Rules we have that the number of ways to form a committee is


"\left(\begin{array}{c}15\\6\end{array}\right)+\left(\begin{array}{c}15\\5\end{array}\right)\left(\begin{array}{c}10\\1\end{array}\right)+\left(\begin{array}{c}15\\4\end{array}\right)\left(\begin{array}{c}10\\2\end{array}\right)="


"=\frac{15!}{6!(9)!}+\frac{15!}{5!(10)!}10+\frac{15!}{4!(11)!}\frac{10!}{2!(8)!}=\frac{15\cdot14\cdot13\cdot12\cdot11\cdot10}{6\cdot5\cdot4\cdot3\cdot2}+
\frac{15\cdot14\cdot13\cdot12\cdot11}{5\cdot4\cdot3\cdot2}10+
\frac{15\cdot14\cdot13\cdot12}{4\cdot3\cdot2}\frac{10\cdot9}{2}="


"=5,005+30,030+61,425=96,460"


Answer: "96,460"